# Weakly Supervised Learning for Facial Behavior Analysis: A Review

### v3

@@ 3.3.1 Expression Recognition
Label Ensembling: Barsoum et al. [64] analyzed four different strategies for effective label ensembling: majority-voting, multi-label learning, probabilistic label drawing, and cross-entropy loss, and showed that the latter two approaches that leverage label distribution outperform the former ones. Jiabei et al. [65] proposed a novel framework of Inconsistent Pseudo Annotations to Latent Truth (IPA2LT) to discover the latent true labels of noisy data with multiple inconsistent annotations. First, predictive models are trained for individually labeled datasets. Next, multiple pseudo-annotations are generated from the trained predictive models for each image, which is used to discover latent true labels based on log-likelihood. Zhang et al. [66] explored the problem of ambiguous and noisy annotations in a decoupled framework using co-training to divide the clean, ambiguous, and noisy samples based on the consistency of the two networks and labels. They further leverage separate regularization strategies for these samples.
Unlike the above-mentioned approaches, Zhang et al. [67] explored a new feature learning perspective to deal with noisy annotations by focusing on the part of features pertinent to noisy annotations. They explored attention consistency by dynamically erasing the input images to prevent the model from overfitting the noisy samples. Li et al. [8] proposed a novel architecture of the Deep Locality-preserving CNN (DLP-CNN) method to obtain more discriminative deep features by preserving the locality closeness while maximizing the inter-class scatters.
3.3.2 Action Unit Recognition
@@ 3.3.3 Discussion
For expression classification, most of the approaches explored deep networks due to the availability of large-scale data with noisy annotations [70, 64, 65]. Relabeling methods [58, 59] depend on the small-loss assumption [71], which may fail to discriminate hard samples and noisy samples as both of them have high loss values during the training process. On the other hand, label assembly methods [64, 65] provide different views of the same sample using different networks, similar to crowdsourcing in real FER applications. They are however computationally expensive, making them less desirable in real-world applications. Label distribution learning is another paradigm [61, 60] allowing us to deal with noisy annotations. Although they offer flexibility in capturing a wide range of emotions, they can lead to spurious distribution depending on the noisy samples. For AU recognition, Zhao et al. [68] explored a clustering approach in the embedding space, while Fabian et al. [69] leveraged the local semantics of AUs that retains the global facial structure for better convergence.
3.4 Experimental Evaluations
3.4.1 Datasets
@@
UNBC-McMaster [72]: The database contains 200 video sequences of 48398 frames captured from 25 participants, who self-identified with shoulder pain. Each frame of the video sequence is labeled with 5 discrete intensity levels of AUs pertinent to pain ($A<B<C<D<E$) obtained by three certified FACS coders. Only the AUs related to pain are considered: AU4, AU6, AU7, AU9, AU10, AU12, AU20, AU25, AU26, AU27 and AU43. However, AU43 is assigned only to binary labels. In addition to the annotations based on FACS, they have also provided labels of discrete pain intensities both at the sequence and frame levels. Prkachin and Solomon Pain Intensity Scale (PSPI) of pain intensities are labeled at frame level with 16 discrete levels from 0-15 and Observer Pain Intensity (OPI) ratings are provided at sequence level on a scale of 0 - 5.
@@
CK+ [73] : The database consists of 593 video sequences captured in controlled laboratory conditions, where the emotions are spontaneously performed by 123 participants. All the video sequences are considered to vary from neutral face to peak formation of the facial expressions and the duration of the sequences varies from 10 to 60 frames. Since the initial labels of these sequences are unreliable, they are further refined based on FACS with seven basic expression labels (including Contempt), and each emotion is defined by a prototypical combination of specific AUs. Out of 593 sequences, it was found that only 327 sequences satisfy the refined labeling strategy, where each video sequence is labeled with the corresponding emotion label. Since the emotion labels are provided at the video level, static approaches assign the emotion label of the video sequence to the last one to three frames that exhibit the peak formation of the expression, and the first frame is considered a neutral frame.
@@
DISFA [75]: Denver Intensity of Spontaneous Facial Action dataset (DISFA) is created using 9 short video clips from YouTube, where the participants of 27 adults are allowed to watch the short video clips pertaining to various emotions. The facial expressions of each of the participants are captured with a high-resolution video of 1024x768 pixels with a frame rate of 20fps resulting in 1,30,754 frames in total. Each of these frames is annotated with AUs along with the discrete intensity levels by FACS expert raters. The AUs related to the expressions in the database are AU1, AU2, AU4, AU5, AU6, AU9, AU12, AU15, AU17, AU20, AU25, and AU26, whose intensities are provided on a six-point ordinal scale (neutral $<$ A $<$ B $<$ C $<$ D $<$ E).
RAF-DB [8]: RAF-DB is a large-scale in-the-wild dataset, which consists of extremely diverse facial images downloaded from the Internet. The obtained images were labeled by 315 annotators, and each image was ensured to be labeled by 40 independent labelers. The final annotations were obtained using crowd-sourcing techniques for seven basic emotions (including neutral expression). The dataset is partitioned into 12,271 training samples and 3,068 test samples. Some of the images are also labeled with compound expressions of 11 classes.
AffectNet [9]: AffectNet is the largest dataset captured in the wild, consisting of 0.4 million images, downloaded from the Internet using three search engines based on expression-related keywords. Its images are labeled with seven basic expressions (excluding neutral), resulting in 2,80,000 samples as training samples and 3,500 images as validation sets.
@@
TABLE II: Performance of state-of-art facial expression classifiers under various WSL settings on most widely evaluated datasets. ’Conventional’ refers to the training and testing partitions provided by the dataset organizers.
@@
TABLE III: Performance of state-of-art AU classifiers under WSL settings on most widely evaluated datasets.
@@ 3.4.2 Experimental Protocol
3.5 Critical Analysis and Results
Facial expressions and AUs evolve over time, therefore facial video may capture significant temporal information about facial behavior. However, temporal information is not well explored in the WSL literature, having only been investigated in a few works [28, 33]. Most of the works on expression detection in the context of incomplete and inaccurate annotations are based on images. Videos have been explored mostly in the context of inexact annotations. Current works on expression detection with inexact annotations used max-pooling or displacement of facial landmarks to extract relevant dynamic information from video sequences. Displacement of facial landmarks was found to capture the temporal dynamics better than max-pooling [30]. [33] is the only work that focused on leveraging the temporal dynamics, showing significant performance improvements over state-of-the-art methods as shown in Table II. Most of these methods for inexact supervision do not leverage the potential of DL models due to limited training data. Few works [37, 41] explored RBMs to capture domain knowledge of expressions and AUs for AU recognition. Most of the works leveraged the domain knowledge of dependencies among AUs and between basic expressions and AUs for AU recognition and achieved comparable performance with each other. In contrast, [41] explored dual learning of AU recognition as the main task and face synthesis as an auxiliary task, showing a significant improvement on several datasets as shown in Table III.
Recently, DL models have been explored for incomplete and inaccurate annotations by leveraging large-scale unlabeled and noisy data sets captured in the wild, respectively. Deep networks have been proven to be robust in handling a wide range of variations such as illumination, pose, identity, etc., better than shallow networks when a large amount of training data is made available [81]. Although pseudo-labeling-based methods [42, 43] and consistency regularization methods [44] have individually shown better performance, effectively leveraging the benefits of both pseudo-labeling and data augmentation was found to be promising, achieving significant improvement over state-of-the-art methods [48] as seen in Figure 5. For AU detection, leveraging expression labels with incomplete AU labels significantly helped to achieve better performance as seen in Fig 6. Specifically, exploring auxiliary tasks seems to be a promising research direction as [41] achieved state-of-the-art performance.
@@ 3.5 Critical Analysis and Results
4 Methods for Regression
In the context of facial behavior analysis, regression deals with the estimation of intensities of facial expressions or AUs. In this section, WSL methods for regression are also categorized into 3 types: inexact, incomplete, and inaccurate annotations. Experimental results obtained with these methods are presented for common evaluation protocols and datasets. Although the problem of regression is not well explored like with classification, it has recently gained much attention due to its immense potential in real-world applications. Few algorithms have been proposed for the task of estimating the intensities of facial expressions or AUs [81], [82] in a fully-supervised setting, the problem is still at a rudimentary stage in the framework of WSL.
4.1 Inexact Annotations
@@
4.1.1 Expression Intensity Estimation
Adria et al. [83] proposed multi-instance dynamic ordinal random fields (MI-DORF) to estimate ordinal intensity levels of frames, where the ordinal variables are modeled as normal distribution and the relationship between the given observation (frame) and latent ordinal value is obtained by projecting the given observation (frame) onto the ordinal line, which is divided by the consecutive overlapping cut-off points of the normal distributions. Praveen et al. [84] further improved performance using the 3D CNN model (I3D in [85]) by integrating MIL into adversarial deep domain adaptation framework [86] for estimating pain intensity, where the source domain is assumed to have fully annotated videos and target domain have periodically annotated weak labels. They have further extended their approach by leveraging the ordinal nature of pain intensity levels and adaptive pooling of instances in the MIL framework [14]. Jianfei et al. [87] proposed an approach for student engagement prediction in the wild using multiple-instance regression, where the input video (bag) was divided into segments (instances), and spatiotemporal features of each segment were fed to an LSTM network.
4.1.2 AU Intensity Estimation
Adria et al. [88] extended the idea of [83] for AU intensity estimation by modeling the relationship between the weak sequence-level label and instance label using two strategies: the maximum or relative values of the instance labels. Unlike the conventional framework of MIL, Zhang et al. [89] explored domain knowledge of relevance using labels of peak and valley frames. Specifically, they have considered three major factors: ordinal relevance, intensity smoothness, and relevance smoothness based on the gradually evolving process of facial behavior.
4.1.3 Discussion
For expression intensity estimation, Adria et al. [83] used features based on facial landmarks and ordinal random fields to model the semantic relationship between the bag label and the latent key instance, as well as temporal dynamics, which has also been extended to AU intensity estimation. Praveen et al. [84, 14] and Jianfei et al. [87] used deep features, while Praveen et al. [84, 14] explored weakly supervised domain adaptation for pain intensity estimation, and Jianfei et al. [87] used LSTM followed by fully connected layers for engagement intensity estimation. For AU intensity estimation, both [88] and [89] used facial landmark-based features and focused on modeling the ordinal relevance of labels along with the temporal smoothness of images in the sequence. However, [88] leverages only the labels of peak frames, whereas [89] uses the labels of both peak and valley frames.
4.2 Incomplete Annotations
@@
4.2.1 Expression Intensity Estimation
@@
4.2.2 AU Intensity Estimation
Yong et al. [91] designed a CNN using annotations of only peak and valley frames, where the parameters of CNN are learned by encoding facial symmetry, temporal intensity ordering, relative appearance similarity, and contrastive appearance difference. Similarly, Zhang et al. [92] explored jointly learning the representation and estimator using only key frames by leveraging temporal label ranking with smoothness as well as feature smoothness. Inspired by the different learning dynamics of multiple AUs, Zhang et al. [93] proposed a context-aware feature and label-fusion modules using an augmented attention mechanism, where feature fusion exploits the spatial relevance of different AUs and label-fusion leverages the partially annotated labels.
Unlike other approaches, Zhang et al. [94] do not rely on keyframes, and extended [91] to joint estimation of the intensities of multiple AUs by introducing a task index to update the corresponding parameters of the fully connected layer. Shangfei et al. [95] also extended the idea of [56] for AU intensity estimation, where RBM is used to model the relationships among AUs, which is explored for AU intensity estimation with partial labels. Sanchez et al. [96] explored self-supervised learning to learn the spatiotemporal representation of large-scale unlabeled videos based on contrastive learning, which is further adapted to randomly labeled datasets without prior information on keyframes.
4.2.3 Discussion
Expression intensity estimation with partial or incomplete annotations is an under-researched problem in the literature. For AU intensity estimation, [91, 94] and [92] explored knowledge-based information such as facial symmetry, temporal intensity ordering, and relative label smoothness, however, [91, 94] used deep features and [92] used facial landmark-based features. [95] explored deep models using RBM to model AU relationships, while [96] and [93] used Resnet-18 to extract features. [93] and [92] are trained on separate models for each AU, whereas [95, 94, 96, 93] leverages jointly training on multiple AUs.

### v4

@@ III-C1 Expression Recognition
Label Ensembling: Barsoum et al. [66] analyzed four different strategies for effective label ensembling: majority-voting, multi-label learning, probabilistic label drawing, and cross-entropy loss, and showed that the latter two approaches that leverage label distribution outperform the former ones. Zeng et al. [67] proposed a novel framework of Inconsistent Pseudo Annotations to Latent Truth (IPA2LT) to discover the latent true labels of noisy data with multiple inconsistent annotations. First, predictive models are trained for individually labeled datasets. Next, multiple pseudo-annotations are generated from the trained predictive models for each image, which is used to discover latent true labels based on log-likelihood. Zhang et al. [68] explored the problem of ambiguous and noisy annotations in a decoupled framework using co-training to divide into clean, ambiguous, and noisy samples based on the consistency of the two networks and labels. They further leverage separate regularization strategies for these samples.
Unlike the above-mentioned approaches, Zhang et al. [69] explored a new feature learning perspective to deal with noisy annotations by focusing on the part of features pertinent to noisy annotations. They explored attention consistency by dynamically erasing the input images to prevent the model from over-fitting the noisy samples. Li et al. [8] proposed a novel architecture of the Deep Locality-preserving CNN (DLP-CNN) method to obtain more discriminative deep features by preserving the locality closeness while maximizing the inter-class scatters.
III-C2 Action Unit Recognition
@@ III-C3 Discussion
For expression classification, most of the approaches explored deep networks due to the availability of large-scale data with noisy annotations [72, 66, 67]. Relabeling methods [60, 61] depend on the small-loss assumption [73], which may fail to discriminate hard (difficult to classify) samples and noisy (mislabeled) samples as both of them have high loss values during the training process. To discriminate the clean samples from the noisy samples, [60] explored ranking regularization whereas [61] showed that modeling the relative uncertainty to different classes is a promising direction as the uncertainty of noisy samples vary with different classes. On the other hand, label ensembling methods [66, 67] provide different views of the same sample using different networks, similar to crowd-sourcing in real FER applications. To estimate the latent true labels of the noisy samples, [67] explored probabilistic modeling of hidden true labels whereas [68] explored co-training based approach to discriminate noisy samples from hard samples. Unlike these two methods, [69] explored a novel direction of refining the noisy samples in the feature space based on the observation that FER models learn noisy samples only on a part of the features instead of whole features. However, label-ensembling methods are computationally expensive, making them less desirable in real-world applications. Label distribution learning is another paradigm [63, 62] allowing us to deal with noisy annotations label distribution of soft labels instead of hard labels. [62] explored multiple auxiliary classifiers to refine the noisy labels in the latent distributions space, whereas [63] showed that label smoothing regularization using soft labels can help in effectively refining the noisy samples. Although they offer flexibility in capturing a wide range of emotions, they can lead to spurious distribution depending on the noisy samples.
For AU recognition, Zhao et al. [70] explored a clustering approach in the embedding space, while Fabian et al. [71] leveraged the local semantics of AUs that retains the global facial structure for better convergence.
III-D Experimental Evaluations
III-D1 Datasets
In this section, we provide some of the widely used datasets for evaluating the WSL approaches for facial behavior analysis.
@@ III-D1 Datasets
UNBC-McMaster [74]: This database contains 200 video sequences of 48398 frames captured from 25 participants, who self-identified with shoulder pain. Each frame of the video sequence is labeled with 5 discrete intensity levels of AUs pertinent to pain ($A<B<C<D<E$) obtained by three certified FACS coders. Only the AUs related to pain are considered: AU4, AU6, AU7, AU9, AU10, AU12, AU20, AU25, AU26, AU27 and AU43. However, AU43 is assigned only to binary labels. In addition to the annotations based on FACS, they have also provided labels of discrete pain intensities both at the sequence and frame levels. Prkachin and Solomon Pain Intensity Scale (PSPI) of pain intensities are labeled at frame level with 16 discrete levels from 0-15 and Observer Pain Intensity (OPI) ratings are provided at sequence level on a scale of 0 - 5.
@@
CK+ [75] : This database consists of 593 video sequences captured in controlled laboratory conditions, where the emotions are spontaneously performed by 123 participants. All the video sequences are considered to vary from neutral face to peak formation of the facial expressions and the duration of the sequences varies from 10 to 60 frames. Since the initial labels of these sequences are unreliable, they are further refined based on FACS with seven basic expression labels (including Contempt), and each emotion is defined by a prototypical combination of specific AUs. Out of 593 sequences, it was found that only 327 sequences satisfy the refined labeling strategy, where each video sequence is labeled with the corresponding emotion label. Since the emotion labels are provided at the video level, static approaches assign the emotion label of the video sequence to the last one to three frames that exhibit the peak formation of the expression, and the first frame is considered a neutral frame.
@@
DISFA [77]: The dataset is created using 9 short video clips from YouTube, where the participants of 27 adults are allowed to watch the short video clips pertaining to various emotions. The facial expressions of each of the participants are captured with a high-resolution video of 1024x768 pixels with a frame rate of 20fps resulting in 1,30,754 frames in total. Each of these frames is annotated with AUs along with the discrete intensity levels by FACS expert raters. The AUs related to the expressions in the database are AU1, AU2, AU4, AU5, AU6, AU9, AU12, AU15, AU17, AU20, AU25, and AU26, whose intensities are provided on a six-point ordinal scale (neutral $<$ A $<$ B $<$ C $<$ D $<$ E).
RAF-DB [8]: This is a large-scale in-the-wild dataset, which consists of extremely diverse facial images downloaded from the Internet. The obtained images were labeled by 315 annotators, and each image was ensured to be labeled by 40 independent labelers. The final annotations were obtained using crowd-sourcing techniques for seven basic emotions (including neutral expression). The dataset is partitioned into 12,271 training samples and 3,068 test samples. Some of the images are also labeled with compound expressions of 11 classes.
AffectNet [9]: This is the largest dataset captured in the wild, consisting of 0.4 million images, downloaded from the Internet using three search engines based on expression-related keywords. Its images are labeled with seven basic expressions (excluding neutral), resulting in 2,80,000 samples as training samples and 3,500 images as validation sets.
@@
TABLE II: Performance of state-of-art facial expression classification approaches under various WSL settings on most widely evaluated datasets. ’Conventional’ refers to the training and testing partitions provided by the dataset organizers. $\dagger$ denotes that the model is trained on both RAF-DB and AffectNet training sets.
@@
TABLE III: Performance of state-of-art AU classification approaches under WSL settings on most widely evaluated datasets.
@@ III-D2 Experimental Protocol
III-E Critical Analysis and Results
Facial expressions and AUs evolve over time, therefore facial video may capture significant temporal information about facial behavior. However, temporal information is less explored in the WSL literature, having only been investigated in a few works [30, 35]. Most of the works on expression detection in the context of incomplete and inaccurate annotations are based on images. Videos have been explored mostly in the context of inexact annotations. Current works on expression detection with inexact annotations used max-pooling or displacement of facial landmarks to extract relevant dynamic information from video sequences. Displacement of facial landmarks was found to capture the temporal dynamics better than max-pooling [32]. [35] is the only work that focused on leveraging the temporal dynamics, showing significant performance improvements over state-of-the-art methods as shown in Table II. Most of these methods for inexact supervision do not leverage the potential of DL models due to limited training data. Few works [39, 43] explored RBMs to capture domain knowledge of expressions and AUs for AU recognition. Most of the works leveraged the domain knowledge of dependencies among AUs and between basic expressions and AUs for AU recognition, achieving comparable performance with each other. In contrast, [43] explored dual learning of AU recognition as the main task and face synthesis as an auxiliary task, showing a significant improvement on several datasets as shown in Table III.
Recently, DL models have been explored for incomplete and inaccurate annotations by leveraging large-scale unlabeled and noisy data sets captured in the wild, respectively. Deep networks have shown robustness in handling a wide range of variations such as illumination, pose, identity, etc., over shallow networks when a large amount of training data is made available [83]. Although pseudo-labeling-based methods [44, 45] and consistency regularization methods [46] have individually shown better performance, effectively leveraging the benefits of both pseudo-labeling and data augmentation was found to be promising, achieving significant improvement over state-of-the-art methods [50] as seen in Figure 5. For AU detection, leveraging expression labels with incomplete AU labels significantly helped to achieve better performance as seen in Fig 6. Specifically, exploring auxiliary tasks seems to be a promising research direction as [43] achieved state-of-the-art performance.
@@ III-E Critical Analysis and Results
IV Methods for Regression
In this section, we present the WSL methods in each scenario along with experimental results and critical analysis for both expressions and AUs similar to that of classification.
IV-A Inexact Annotations
@@
IV-A1 Expression Intensity Estimation
Ruiz et al. [84] proposed multi-instance dynamic ordinal random fields (MI-DORF) to estimate ordinal intensity levels of frames, where the ordinal variables are modeled as normal distribution and the relationships between the given observation (frame) and latent ordinal value is obtained by projecting the given observation (frame) onto the ordinal line, which is divided by the consecutive overlapping cut-off points of the normal distributions. Praveen et al. [85] further improved performance using the 3D CNN model (I3D in [86]) by integrating MIL into adversarial deep domain adaptation framework [87] for estimating pain intensity, where the source domain is assumed to have fully annotated videos and target domain have periodically annotated weak labels. They have further extended their approach by leveraging the ordinal nature of pain intensity levels and adaptive pooling of instances in the MIL framework [14].
IV-A2 AU Intensity Estimation
Ruiz et al. [88] extended the idea of [84] for AU intensity estimation by modeling the relationships between the weak sequence-level label and instance label using two strategies: the maximum or relative values of the instance labels. Unlike the conventional framework of MIL, Zhang et al. [89] explored domain knowledge of relevance using labels of peak and valley frames. Specifically, they have considered three major factors: ordinal relevance, intensity smoothness, and relevance smoothness based on the gradually evolving process of facial behavior.
IV-A3 Discussion
For expression intensity estimation, Ruiz et al. [84] used features based on facial landmarks and ordinal random fields to model the semantic relationships between the bag label and the latent key instance, as well as temporal dynamics, which has also been extended to AU intensity estimation. Praveen et al. [85, 14] explored deep models with weakly supervised domain adaptation for pain intensity estimation, where the source domain is exploited to deal with limited representative data.
For AU intensity estimation, both [88] and [89] used facial landmark-based features and focused on modeling the ordinal relevance of labels along with the temporal smoothness of images in the sequence. However, [88] leverages only the labels of peak frames, whereas [89] uses the labels of both peak and valley frames. Exploiting both peak and valley frames seems to be efficient in relatively modeling the AU intensity levels instead of relying only on peak frames. The problem of expression and AU intensity estimation with inexact annotations is still at rudimentary level, though it has been well explored in fully supervised setting [83, 90].
IV-B Incomplete Annotations
@@
IV-B1 Expression Intensity Estimation
@@
IV-B2 AU Intensity Estimation
Zhang et al. [92] designed a CNN using annotations of only peak and valley frames, where the parameters of the CNN are learned by exploiting domain knowledge of facial symmetry, temporal intensity ordering, relative appearance similarity, and contrastive appearance difference of unlabeled frames. In addition to the domain knowledge of unlabeled frames, [93] explored joint learning of feature representations and estimator of AUs, where facial landmark based features are transformed to a latent space, which is learnt along with the AU estimators. Motivated by different learning dynamics of multiple AUs, Zhang et al. [94] proposed a context-aware feature and label-fusion modules using an augmented attention mechanism, where feature fusion exploits the spatial relevance of different AUs and label-fusion leverages the partially annotated labels to capture the temporal dynamics of AUs.
Unlike other approaches, Zhang et al. [95] do not rely on keyframes, and extended [92] to joint estimation of the intensities of multiple AUs by introducing a task index to update the corresponding parameters of the fully connected layer. Wang et al. [96] also extended the idea of [58] for AU intensity estimation, where RBM is used to model the relationships among AUs, which is explored for AU intensity estimation with partial labels. Sanchez et al. [97] explored self-supervised learning to learn the spatiotemporal representation of large-scale unlabeled videos based on contrastive learning, which is further adapted to partially (randomly) labeled datasets without prior information of key-frames.
IV-B3 Discussion
Expression intensity estimation with partial or incomplete annotations is an under-researched problem in the literature. For AU intensity estimation, [92, 95] and [93] explored knowledge-based information such as facial symmetry, temporal intensity ordering, and relative label smoothness. However, [92, 95] used deep features with labels of only peak and valley frames (1% labeled frames), whereas [93] used facial landmark-based features with additional randomly selected labeled frames (total 6% labeled frames). Unlike these two approaches, [94] showed that different AUs has different learning dynamics, and explored context-aware feature fusion for learning feature representations of AUs. [96] explored deep models using RBM to model AU relationships, while [97] explored self-supervised learning framework for learning spatiotemporal features with pretraining on large-scale dataset. Due to the availability of large-scale unlabeled frames, most of these approaches except [93] explored DL models . [94] and [93] are trained on separate models for each AU, whereas [96, 95, 97, 94] leverages jointly training on multiple AUs. Considering the differences in learning dynamics of different AUs seems to be promising line of research for better performance.